OpenStudy (anonymous):

come up with an expression equivalent to 1-tan^4(x)

OpenStudy (anonymous):

the answer would look similar to sec-sec or something like that. All I know is that sec is in the answer

OpenStudy (vishweshshrimali5):

See.. first start with using the identity: \(a^2 - b^2 = (a-b)(a+b)\)

OpenStudy (anonymous):

so (1-tan^2)(1+tan^2) ???

OpenStudy (vishweshshrimali5):

Great!!

OpenStudy (vishweshshrimali5):

Now what is 1+ tan^2 ?

OpenStudy (anonymous):

sec^2x

OpenStudy (vishweshshrimali5):

Good!

OpenStudy (vishweshshrimali5):

Anything you can think about our second buddy... 1 - tan^2 x ?

OpenStudy (anonymous):

? would I have to separate that?

OpenStudy (anonymous):

(1-tanx)(1+tanx) ???

OpenStudy (vishweshshrimali5):

I think just leave that 1-tan^2 x

OpenStudy (anonymous):

oh.

OpenStudy (vishweshshrimali5):

You wanted a sec^2 right? You got that already :)

OpenStudy (anonymous):

I have it. but the answer choices have sec on both sides

OpenStudy (anonymous):

answer choices r faded so its hard to see

OpenStudy (vishweshshrimali5):

Ohh well that's a tough task then...

OpenStudy (anonymous):

like one is 2sec^2x-sec^4x......or -2sec^2x-sec^4x

OpenStudy (anonymous):

thats 2 answer choices right there

OpenStudy (vishweshshrimali5):

hmm... well let's see what we have till now... \(1-\tan^4 x\) \(=(1-\tan^2 x)(\sec^2x)\)

OpenStudy (vishweshshrimali5):

There are several things we can do for 1 - tan^2 Like we have.. \[\cos(2x) = \frac{1- \tan^2 x}{1+\tan^2 x}\]

OpenStudy (vishweshshrimali5):

So, if I replace 1- tan^2 x with cos(2x) (sec^2x)

OpenStudy (anonymous):

for example.... Which of the following expressions is equal to 1-sec^4x???? A. -2tan^2x-tan^4x C. 2tan^2x-tan^4x B. tan^2x-tan^4x D. -2tan^2x+tan^4x The answer was A..... So for the problem we were doing I thought it would be similar.

OpenStudy (vishweshshrimali5):

I don't think there is any use of simplifying it any further... let's just go with cos(2x)*sec^4 x

OpenStudy (anonymous):

Same here, but its not any one of the answer choices.

OpenStudy (anonymous):

But I appreciate your help!

OpenStudy (vishweshshrimali5):

Your welcome :)

OpenStudy (faiqraees):

So what are your choices?

OpenStudy (faiqraees):

Btw you can write 1-tan²x as 2 - sec²x

OpenStudy (anonymous):

Very hard to see as i did get a faded print out but from what I can manage to make out I have A.2sec^4x+sec^4x B. sec^2x-sec^4x C. 2sec^2x-sec^4x D. -2sec^2x-sec^4x

OpenStudy (faiqraees):

yeah so C is correct

OpenStudy (anonymous):

so my answer would be positive?

OpenStudy (faiqraees):

(2-sec²x)sec²x 2sec²x-sec^4x

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

thanks a lot